# Learning Harmonized Representations for Speculative Sampling

### v2

@@ Abstract
Speculative sampling is a promising approach to accelerate the decoding stage for Large Language Models (LLMs). Recent advancements that leverage target LLM’s contextual information, such as hidden states and KV cache, have shown significant practical improvements. However, these approaches suffer from inconsistent context between training and decoding. We also observe another discrepancy between the training and decoding objectives in existing speculative sampling methods. In this work, we propose a solution named HArmonized Speculative Sampling (HASS) that learns harmonized representations to address these issues. HASS accelerates the decoding stage without adding inference overhead through harmonized objective distillation and harmonized context alignment. Experiments on four LLaMA models demonstrate that HASS achieves 2.81x-4.05x wall-clock time speedup ratio averaging across three datasets, surpassing EAGLE-2 by 8%-20%.
@@ 1 Introduction
Another discrepancy is also observed between the objectives of the training and decoding stages. During the decoding stage, the objective of the draft model is to propose tokens that the target LLM is likely to assign high probabilities to (Li et al., 2024c; Miao et al., 2024; Sun et al., 2024). In this scenario, the draft model should focus more on recalling the desired tokens, while the specific order of these tokens can be somewhat de-emphasized. Moreover, most LLM applications perform nucleus sampling (Holtzman et al., 2019) or top-k sampling (Fan et al., 2018). For these decoding objectives, tokens with high probabilities play a more significant role in determining the output. Therefore, to develop efficient draft models, their training objectives should consider these properties encountered in the decoding stage. To the best of our knowledge, previous works on training draft models for speculative sampling have largely overlooked these decoding considerations.
Figure 1: Speedup ratios of different methods on LLaMA2-Chat 7B/13B and LLaMA3-Instruct 8/70B with temperature T $\in\{0,1\}$, averaging over MT-bench, HumanEval, and GSM8K datasets.
Figure 2: We use EAGLE (Li et al., 2024b) as an example to illustrate the context misalignment, where the speculation starts from timestep $t$. $f^{(l)}$ and $f^{(s)}$ represent hidden states from the target LLM and the draft model, respectively. When decoding draft token $x_{t+2}$, the input context is inconsistent between training and decoding.
@@
We conduct experiments across dialogue, code generation, and mathematical reasoning using the MT-bench, HumanEval, and GSM8K datasets, respectively. Building with EAGLE-2 (Li et al., 2024c), HASS achieves 8%-16% acceptance length improvement over it on LLaMA2-Chat 7/13B and LLaMA3-Instruct 8/70B, resulting in 2.81x-4.05x wall-clock time acceleration compared with the vanilla inference on NVIDIA H800 GPU.
@@ 2 Preliminary
Leviathan et al. (2023) demonstrate that the wall-clock time improvement ratio is directly proportional to $\tau$, while the arithmetic operation increment ratio is inversely proportional to $\tau$. Consequently, $\tau$, also known as the acceptance length, plays a crucial role in determining the performance of acceleration. This analysis also applies when using multiple draft sequences (Miao et al., 2024; Li et al., 2024c; b; Sun et al., 2024). Note that $\tau$ is closely related to the distribution gap between the target LLM and the draft model. With efficient decoding requirements, the draft model typically has limited capacity, resulting in a significant distribution gap compared to the target LLM. Fortunately, during inference, the acceptance rate is primarily influenced by the alignment of distributions on the desired tokens, i.e., the tokens to which the target LLM assigns high probabilities. However, previous speculative sampling works mainly focus on the entire vocabulary set w.r.t. knowledge distillation from the target LLM (Li et al., 2024b; Zhou et al., 2023), thereby disconnecting the training process from the practical decoding requirements.
EAGLE (Li et al., 2024b) is a lightweight draft model design, as shown in Figure 2. During decoding, it utilizes the LM Head of the target LLM to generate draft tokens. Specifically, we assume that the speculation starts from timestep $t$, meaning the first draft token is at timestep $t+1$. To generate the draft token $x_{t+1}$, the target LLM’s hidden state $f_{t-1}^{(l)}$ in the second-to-top layer is concatenated with the embedding $e_{t}$ to perform the input of the draft model. During training, EAGLE constructs a regression task between $f^{(l)}$s and the predicted hidden states $f^{(s)}$s of the draft model. However, due to the auto-regressive decoding, the draft model only accesses the target LLM’s features at the beginning of the speculation. It uses the features produced by the draft models as input for subsequent steps. This context misalignment, stemming from feature inaccuracies, leads to error accumulation and hinders the performance of generating later draft tokens (Li et al., 2024b; Du et al., 2024). EAGLE-2 (Li et al., 2024c) employs the same model design but works on dynamic drafting structures instead of a static tree structure during the decoding stage, yet the aforementioned issue remains unresolved.
@@ 3.2 Harmonized Context Alignment
HASS follows a context alignment schema that aligns training and decoding on their contexts. In particular, HASS divides the training procedure into $n$ steps, enabling the draft model to utilize contextual features consistent with those in the decoding stage. The procedure is outlined as follows:
  • The first step mirrors the training stage of EAGLE. At timestep $t+1$, the draft model takes the target model’s feature $f^{(l)}_{t}$ as input and produces the draft feature $f^{(s_{1})}_{t+1}$. In this step, the attention mask remains as the original causal mask without any modification.
  • In the second step, features from the first step are incorporated. For instance, in the self-attention mechanism at timestep $t+1$, $f^{(s_{1})}_{t}$ is used to calculate the current query. Keys and values are derived from $f^{(l)}_{:t}\oplus f^{(s_{1})}_{t}$, where $\oplus$ denotes concatenation and $f^{(l)}_{:t}$ includes features from timesteps earlier than $t$. The attention mask is adjusted to ensure that the previous feature seen by $f^{(s_{1})}_{i}$ is always $f^{(l)}_{i-1}$, as shown in the ‘HASS Training Step 2’ part of Figure 3.
@@
It takes $n$ times the training overhead of EAGLE while maintaining the same decoding overhead. We empirically demonstrate that the acceleration effect converges with a small $n$ so that the training of HASS is cost-efficient.
@@
Table 1: Acceptance lengths $\tau$ of different methods on MT-bench, HumanEval, and GSM8K datasets with temperature $\text{T}\in\{0,1\}$. L2 represents LLaMA2-Chat, while L3 represents LLaMA3-Instruct.
@@
Table 2: Speedup ratios of different methods on MT-bench, HumanEval, and GSM8K datasets with temperature T $\in\{0,1\}$. L2 represents LLaMA2-Chat, while L3 represents LLaMA3-Instruct.
@@ 4.1 Experimental Setup
Target LLMs. LLaMA2-Chat 7B/13B and LLaMA3-Instruct 8/70B.
@@
Comparisons. The vanilla auto-regressive decoding is taken as the baseline, which serves as the benchmark for speedup ratios (1.00x). We compare HASS with recent lossless speculative sampling methods, including PLD (Saxena, 2023), Lookahead (Fu et al., 2023), EAGLE (Li et al., 2024b), and EAGLE-2 (Li et al., 2024c).
Implementation. Our code is built based on EAGLE-2’s open-source repository [^1] https://github.com/SafeAILab/EAGLE. Experiments on EAGLE and EAGLE-2 reuse draft model weights trained by Li et al. (2024b). For harmonized objective distillation, K is set as 10, and the loss of harmonized objective distillation is added to EAGLE’s original loss with a coefficient of $w=1.0$. For harmonized context alignment, the draft model is aligned for 3 steps during training. For dynamic tree structure, we set the total number of draft tokens to 60 for all experiments with a draft tree depth of 6. We keep other settings, such as the fixed training dataset and the optimizer, consistent with EAGLE-2.
@@ 4.2 Effectiveness & Ablation Study
In this section, we first evaluate the effectiveness of HASS by comparing it with existing speculative sampling methods on acceptance length and speedup ratio. Then, we conduct ablation studies on harmonized objective distillation and harmonized context alignment.
@@ 4.2.1 Effectiveness
We present different methods’ acceptance lengths and speedup ratios across three datasets in Tables 1 and 2, respectively. HASS performs the largest acceptance length and highest speedup ratio across all datasets and LLMs we tested. Most methods achieve their best performance on the HumanEval dataset, as the fixed templates in the code generation task are easier to draft and accelerate. Though PLD and Lookahead are free of training, they consistently show poorer performance than EAGLE, EAGLE-2, and HASS.
@@ 4.2.3 Ablation Study on Harmonized Context Alignment
As the first training step of HASS is the same as EAGLE-2, we continually train EAGLE-2’s draft model weights with the Top-K loss and consider it the baseline. Without harmonized context alignment (EAGLE-2 + Top-K), the draft model performs the worst across all datasets. Training with 3/4 steps of harmonized context alignment generally obtains the most considerable acceptance length. When training with five steps of context alignment, the acceptance length decreases. We believe this is caused by the draft model’s limited capacity, as it predicts less accurately on former steps’ tokens when paying too much attention to the latter ones. Figure 5 shows the acceptance rate $\alpha$ across speculation steps on the MT-bench dataset following Li et al. (2024c). In later speculation steps, HASS performs better acceptance rates than EAGLE-2, demonstrating the effectiveness of harmonized context alignment.
Figure 5: Acceptance rates $\alpha$ of HASS and EAGLE-2 across different speculation steps on the MT-bench dataset with temperature T $\in\{0,1\}$.
@@
Figure 6: Acceptance rates $\alpha$ of HASS with different reweight factors $\beta$ for harmonized context alignment. The results are conducted on LLaMA3-Instruct 70B over the MT-bench dataset with temperature T $\in\{0,1\}$.
@@ 6 Conclusion
Appendix for ‘Learning Harmonized Representations for Speculative Sampling’
Appendix A Implementation of Harmonized Context Alignment
@@
Appendix B Harmonized Context Alignment on Tokens
@@
Table 6: Acceptance lengths $\tau$ of applying feature and token alignment to EAGLE-2’s draft model weights, where ‘Token ($x$)’ denotes tokens from training data are replaced by draft-model-generated tokens with a probability of $x$. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Figure 7: Acceptance rates $\alpha$ of applying feature and token alignment to EAGLE-2’s draft model weights, where ‘Token ($x$)’ denotes tokens from training data are replaced by draft-model-generated tokens with a probability of $x$. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Feature only alignment brings the best performance, while token alignment leads to degeneration. With the probability of applying token alignment increasing from 0.1 to 1.0, the acceptance length decreases consistently. As shown in Figure 7, more token alignment generally causes lower acceptance rates. As a result, training with tokens generated by the draft model in harmonized context alignment hurts the acceleration performance.
Appendix C Hyper-Parameters of Top-K Loss
@@
Appendix D Self-Distillation
Following existing speculative sampling methods (Zhou et al., 2023; Cai et al., 2024), we further use target-model-generated outputs to distill the draft model from the target model’s real output distribution, dubbed as self-distillation. In specific, we feed the prompts from the ShareGPT dataset into the target models recursively and collect the responses as multi-turn conversations for self-distillation. To study the effect of self-distillation, we conduct experiments by training the draft model with fixed data or model-generated data and summarize the results in Table 8.

### v3

@@ Abstract
Speculative sampling is a promising approach to accelerate the decoding stage for Large Language Models (LLMs). Recent advancements that leverage target LLM’s contextual information, such as hidden states and KV cache, have shown significant practical improvements. However, these approaches suffer from inconsistent context between training and decoding. We also observe another discrepancy between the training and decoding objectives in existing speculative sampling methods. In this work, we propose a solution named HArmonized Speculative Sampling (HASS) that learns harmonized representations to address these issues. HASS accelerates the decoding stage without adding inference overhead through harmonized objective distillation and harmonized context alignment. Experiments on four LLaMA models demonstrate that HASS achieves 2.81x-4.05x wall-clock time speedup ratio averaging across three datasets, surpassing EAGLE-2 by 8%-20%. The code is available at https://github.com/HArmonizedSS/HASS.
@@ 1 Introduction
Another discrepancy is also observed between the objectives of the training and decoding stages. During the decoding stage, the objective of the draft model is to propose tokens that the target LLM is likely to assign high probabilities to (Li et al., 2024c; Miao et al., 2024; Sun et al., 2024). In this scenario, the draft model should focus more on recalling the desired tokens, while the specific order of these tokens can be somewhat de-emphasized. Moreover, most LLM applications perform nucleus sampling (Holtzman et al., 2020) or top-k sampling (Fan et al., 2018). For these decoding objectives, tokens with high probabilities play a more significant role in determining the output. Therefore, to develop efficient draft models, their training objectives should consider these properties encountered in the decoding stage. To the best of our knowledge, previous works on training draft models for speculative sampling have largely overlooked these decoding considerations.
Figure 1: Speedup ratios of different methods on LLaMA2-Chat 7/13B and LLaMA3-Instruct 8/70B with temperature T $\in\{0,1\}$, averaging over MT-bench, HumanEval, and GSM8K datasets.
Figure 2: We use EAGLE (Li et al., 2024b) as an example to illustrate the context misalignment, where the speculation starts from timestep $t$. $f^{(l)}$ and $f^{(s)}$ represent hidden states from the target LLM and the draft model. When decoding draft token $x_{t+2}$, the input context is inconsistent between training and decoding.
@@
We conduct experiments across dialogue, code generation, and mathematical reasoning tasks using the MT-bench, HumanEval, and GSM8K datasets, respectively. Building with EAGLE-2 (Li et al., 2024c), HASS achieves 8%-16% acceptance length improvement over it on LLaMA2-Chat 7/13B and LLaMA3-Instruct 8/70B, resulting in 2.81x-4.05x wall-clock time acceleration compared with the vanilla inference on NVIDIA H800 GPU.
@@ 2 Preliminary
Leviathan et al. (2023) demonstrate that the wall-clock time improvement ratio is directly proportional to $\tau$, while the arithmetic operation increment ratio is inversely proportional to $\tau$. Consequently, $\tau$, also known as the acceptance length, plays a crucial role in determining the performance of acceleration. This analysis also applies when using multiple draft sequences (Miao et al., 2024; Li et al., 2024c; b; Sun et al., 2024). Note that $\tau$ is closely related to the distribution gap between the target LLM and the draft model. With efficient decoding requirements, the draft model typically has limited capacity, resulting in a significant distribution gap compared to the target LLM. Fortunately, during inference, the acceptance rate is primarily influenced by the alignment of distributions on the desired tokens, i.e., the tokens to which the target LLM assigns high probabilities. However, previous speculative sampling works mainly focus on the entire vocabulary set w.r.t. knowledge distillation from the target LLM (Li et al., 2024b; Zhou et al., 2024), thereby disconnecting the training process from the practical decoding requirements.
EAGLE (Li et al., 2024b) is a lightweight draft model design, as shown in Figure 2. During decoding, it utilizes the LM Head of the target LLM to generate draft tokens. Specifically, we assume that the speculation starts from timestep $t$, meaning the first draft token is at timestep $t+1$. To generate the draft token $x_{t+1}$, the target LLM’s hidden state $f_{t-1}^{(l)}$ in the second-to-top layer is concatenated with the embedding $e_{t}$ to perform the input of the draft model. During training, EAGLE constructs a regression task between $f^{(l)}$s and the predicted hidden states $f^{(s)}$s of the draft model. However, due to the auto-regressive decoding, the draft model only accesses the target LLM’s features at the beginning of the speculation. It uses the features produced by itself as input for subsequent steps. This context misalignment, stemming from feature inaccuracies, leads to error accumulation and hinders the performance of generating later draft tokens (Li et al., 2024b; Du et al., 2024). EAGLE-2 (Li et al., 2024c) employs the same model design but works on dynamic drafting structures instead of a static tree structure during the decoding stage, yet the aforementioned issue remains unresolved.
@@ 3.2 Harmonized Context Alignment
HASS follows a context alignment schema that aligns training and decoding on their contexts. The training procedure of HASS is divided into $n$ steps, enabling the draft model to utilize contextual features consistent with those in the decoding stage and addressing the context inconsistency by adapting the inaccurate features generated in previous HASS training steps. Specifically, it is achieved by first taking the inaccurate feature from the last draft model as query, and then considering the inaccuracy accumulation in key-value part of the transformer block.
Formally, in the HASS training step $j$, given input token sequence $x_{1},x_{2},..,x_{T}$, we optimize the draft model $\mathcal{M}^{(s)}$ with the following objective function:
[…]
$P^{(s)}(x_{t+1}|x_{1},\dots,x_{t})$ $=\text{Head}(f_{t+1}^{(s_{j})})$
$=\text{Head}(\mathcal{M}^{(s)}(\overbrace{\underbrace{f_{t}^{(s_{j-1})}}_{\text{query}}}^{\text{From last draft}},\underbrace{\overbrace{f_{1}^{(l)}\oplus\cdots\oplus f_{t-j+1}^{(l)}}^{\text{From target LLM}}\oplus\overbrace{f_{t-j+2}^{(s_{1})}\oplus\cdots\oplus f_{t}^{(s_{j-1})}}^{\text{From previous draft models}}}_{\text{key \& value}})),$
$P^{(l)}$ is the auto-regressive probability distribution provided by the target LLM, the Aux-loss consists of the proposed Top-K loss and the feature regression loss (following EAGLE), Head and $\oplus$ stand for the language modeling head and the concatenation operation respectively. When training tokens in the entire sequence in parallel, the above formulation adapts the inaccurate features from previous $j-1$ steps for all positions except the first $j-1$ positions. Note that compared to EAGLE, HASS takes additional training overhead due to the extra $n-1$ training steps for adapting inaccurate features, while maintaining the same decoding overhead. To accelerate the training procedure, we propose a modification to the attention mask mechanism, as outlined below:
  • The first step mirrors the training stage of EAGLE. At timestep $t+1$, the draft model takes the target model’s feature $f^{(l)}_{t}$ as input and produces the draft feature $f^{(s_{1})}_{t+1}$. In this step, the attention mask remains the same as the original causal mask without any modification.
  • In the second step, features from the first step are incorporated. For instance, in the self-attention mechanism at timestep $t+1$, $f^{(s_{1})}_{t}$ is used to calculate the current query. Keys and values are derived from $f^{(l)}_{:t}\oplus f^{(s_{1})}_{t}$, where $f^{(l)}_{:t}$ includes features from timesteps earlier than $t$. The attention mask is adjusted to ensure that the previous feature seen by $f^{(s_{1})}_{i}$ is always $f^{(l)}_{i-1}$, as shown in the ‘HASS Training Step 2’ part of Figure 3.
@@
We empirically demonstrate that the acceleration effect converges with a small $n$ so that the training of HASS is cost-efficient. The actual training overhead of HASS in terms of training speed, computational cost, and GPU memory is investigated in Appendix A.8.
@@
Table 1: Acceptance lengths $\tau$ of different methods on MT-bench, HumanEval, and GSM8K datasets with temperature $\text{T}\in\{0,1\}$. L2 represents LLaMA2-Chat, while L3 represents LLaMA3-Instruct. SpS stands for Vanilla Speculative Sampling, while V-68M and L-68M represent Vicuna-68M and LLaMA-68M, which are the draft models of SpS.
@@
Table 2: Speedup ratios of different methods on MT-bench, HumanEval, and GSM8K datasets with temperature T $\in\{0,1\}$. L2 represents LLaMA2-Chat, while L3 represents LLaMA3-Instruct. SpS stands for Vanilla Speculative Sampling, while V-68M and L-68M represent Vicuna-68M and LLaMA-68M, which are the draft models of SpS.
@@ 4.1 Experimental Setup
Target LLMs. LLaMA2-Chat 7/13B and LLaMA3-Instruct 8/70B.
@@
Comparisons. The vanilla auto-regressive decoding is taken as the baseline, which serves as the benchmark for speedup ratios (1.00x). We compare HASS with recent lossless speculative sampling methods, including PLD (Saxena, 2023), Lookahead (Fu et al., 2023), Vanilla Speculative Sampling (Chen et al., 2023), Medusa (Cai et al., 2024), EAGLE (Li et al., 2024b), and EAGLE-2 (Li et al., 2024c). PLD and Lookahead are free of traning, which respectively use string matched from the prompt and cached n-grams as draft tokens instead of generating draft tokens from a draft model’s predicted probability distribution. Therefore, the results of PLD and Lookahead under temperature = 1 are not reported in Table 1.
Implementation. Our code is built based on EAGLE-2’s open-source repository [^1] https://github.com/SafeAILab/EAGLE. Experiments on EAGLE and EAGLE-2 reuse draft model weights trained by Li et al. (2024b). For harmonized objective distillation, K is set as 10, and the loss of harmonized objective distillation is added to EAGLE’s original loss with a coefficient of $w=1.0$. For harmonized context alignment, the draft model is aligned for 3 steps during training. For dynamic tree structure, we set the total number of draft tokens to 60 for all experiments with a draft tree depth of 6. We keep other settings, such as the fixed training dataset, i.e., the ShareGPT [^2] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered dataset with 68,000 dialogues, and the optimizer, consistent with EAGLE-2.
@@ 4.2 Effectiveness & Ablation Study
In this section, we first evaluate the effectiveness of HASS by comparing it with existing speculative sampling methods on acceptance length and speedup ratio. Then, we conduct ablation studies on harmonized objective distillation and harmonized context alignment. Inspired by Yi et al. (2024), we further conduct experiments by training on different proportions of the ShareGPT dataset to investigate HASS’s scalability in the face of data sparsity (see Appendix A.6), and by evaluating on the translation tasks to investigate HASS’s robustness across different task types (see Appendix A.7). As shown from the results, HASS is more scalable than EAGLE-2 with fewer training data and achieves promising improvements over EAGLE-2 on translation tasks consistent with results on MT-bench, HumanEval, and GSM8K.
@@ 4.2.1 Effectiveness
We present different methods’ acceptance lengths and speedup ratios across three datasets in Tables 1 and 2, respectively. HASS performs the largest acceptance length and highest speedup ratio across all datasets and LLMs we tested. Most methods achieve their best performance on the HumanEval dataset, as the fixed templates in the code generation task are easier to draft and accelerate. Though PLD and Lookahead are free of training, they consistently show poorer performance than Medusa, EAGLE, EAGLE-2, and HASS.
@@ 4.2.3 Ablation Study on Harmonized Context Alignment
As the first training step of HASS is the same as EAGLE-2, we continually train EAGLE-2’s draft model weights with the Top-K loss and consider it the baseline. Without harmonized context alignment (EAGLE-2 + Top-K), the draft model performs the worst across all datasets. Training with 3/4 steps of harmonized context alignment generally obtains the most considerable acceptance length. When training with 5 steps of context alignment, the acceptance length decreases. We believe this is caused by the draft model’s limited capacity, as it predicts less accurately on former steps’ tokens when paying too much attention to the latter ones. Figure 5 shows the acceptance rate $\alpha$ across speculation steps on the MT-bench dataset following Li et al. (2024c). In later speculation steps, HASS performs better acceptance rates than EAGLE-2, demonstrating the effectiveness of harmonized context alignment.
Figure 5: Acceptance rates $\alpha$ (%) of HASS and EAGLE-2 across different speculation steps on the MT-bench dataset with temperature T $\in\{0,1\}$.
@@
Figure 6: Acceptance rates $\alpha$ (%) of HASS with different reweight factors $\beta$ for harmonized context alignment. The results are conducted on LLaMA3-Instruct 70B over the MT-bench dataset with temperature T $\in\{0,1\}$.
@@ 6 Conclusion
Appendix A Appendix
A.1 Implementation of Harmonized Context Alignment
@@
A.2 Harmonized Context Alignment on Tokens
@@
Table 6: Acceptance lengths $\tau$ of applying feature and token alignment to EAGLE-2’s draft model weights, where ‘Token ($x$)’ denotes tokens from training data being replaced by draft-model-generated tokens with a probability of $x$. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Figure 7: Acceptance rates $\alpha$ (%) of applying feature and token alignment to EAGLE-2’s draft model weights, where ‘Token ($x$)’ denotes tokens from training data being replaced by draft-model-generated tokens with a probability of $x$. The results are conducted on LLaMA2-Chat 7B over the MT-bench dataset with temperature T $\in\{0,1\}$.
Feature only alignment brings the best performance, while adding token alignment leads to degeneration. With the probability of applying token alignment increasing from 0.1 to 1.0, the acceptance length decreases consistently. As shown in Figure 7, more token alignment generally causes lower acceptance rates. As a result, training with tokens generated by the draft model in harmonized context alignment hurts the acceleration performance.
A.3 Hyper-Parameters of Top-K Loss
@@
A.4 Self-Distillation
In the main text, we use the fixed ShareGPT dataset to train draft models for a fair comparison with EAGLE and EAGLE-2. Following existing speculative sampling methods (Zhou et al., 2024; Cai et al., 2024), we further use target-model-generated outputs to distill the draft model from the target model’s real output distribution, dubbed as self-distillation. In specific, we feed the prompts from the ShareGPT dataset into the target models recursively with temperature set to 0 and collect the responses as multi-turn conversations for self-distillation following Li et al. (2024b). To study the effect of self-distillation, we conduct experiments on HASS and EAGLE-2 by training the draft model with fixed data or model-generated data and summarize the results in Table 8.